If the earth has a circumfrence of 60,000Km, how fast must a satelite be traveling if it is 300Km from the surface to remain in orbit?
Hint: gravity is 9.8m/s2
1
Expert's answer
2012-09-28T07:53:12-0400
mg=R*m*w^2
sqrt(g/R)=w (angular speed)

R=h+(60000/2*pi)

w=sqrt(9.8/(h+60000/2*pi))

w=0.33 (angular speed)